OpenStudy (anonymous):

A disk initially at rest is given an angular acceleration of 12.0 rad/ s^2• What is its angular velocity after 10 s? What is its frequency then in revolutions per minute (rev/m in)? How many revolutions does it make during this time?

OpenStudy (irishboy123):

A) dΩ = 12 * 10 = 120rad/s B) 120rad/s = 120*60 = 7,200rad/min = 1,146 rpm (because 2π rad = 1 rev) C) average rpm (because of linear angular acceleration) = 573rpm. t = 1/6mins, revs = 95.5